54.) simplify...(x+2)(x²+x+5)...just need help with this one. please
Answer by stanbon(75887)

simplify...(x+2)(x²+x+5)
= x(x^2+x+5) + 2(x^2+x+5)
= x^3 + x^2 + 5x + 2x^2 + 2x + 10
= x^3 + 3x^2 + 7x + 10
===========================
Cheers,
Stan H.